OpenStudy (anonymous):

z1= sqrt2-sqrt 2i, z2 = 1 − i Write z1 and z2in polar form. (Express θ in radians. Let 0 ≤ θ < 2π.)Find the product z1z2and the quotients z1/z2 and 1/z1 . (Express your answers in polar form with θ expressed in radians.)

OpenStudy (danjs):

Each polar point is a distance and an angle remember to relate cartesian rectangular coord to the polar , use a right triangl the x coord would be distance times cos(angle) y coord will be the distance times cos(angle)

OpenStudy (danjs):

x = r*cos(angle) y = r* sin(angle)

OpenStudy (danjs):

the angle that line makes with the +x axis is the same as the tangent of the ratio of y/x tan(angle)=y/x

OpenStudy (danjs):

you can write complex numbers in polar form the same as usual z = a + b*i

OpenStudy (danjs):

a is x coordinate b is y coord